OpenStudy (nthenic_oftime):

Find the vertical asymptotes, if any, of the graph of the rational function. f(x) = x/x^2+1 so for this one id set x^2+1 = 0 and the x=1? @jhannybean A. X=1 B. X=-1 C.X=-1, X=1 D. no vertical asymptotes its A correct?

OpenStudy (jhannybean):

\[x^2+1=0 \qquad x^2 = -1 \implies \text{no asymptotes}\] There are no real values that satisfy x

OpenStudy (jhannybean):

no vertical asymptotes*

OpenStudy (jhannybean):

\\[x^2=-1 \qquad x = \pm \sqrt{-1} = \pm i\]

OpenStudy (nthenic_oftime):

:)